Compute for the molar solubility of cadmium hydroxide which is slightly soluble in water if the Ksp value is 6.0 x 10-12.


1
Expert's answer
2021-03-01T02:35:07-0500

From the balanced net ionic equation;

"Cd^{2+}_{(aq)}+2OH^{-}_{(aq)}\leftrightharpoons Cd(OH)_{2_{(aq)}}"

"K_{sp}=[Cd^{2+}][OH^-]"

"6.0\times 10^{-12}=(x).(2x^2)"

"4x^3=6.0\times 10^{-12}"

"x^3=1.5\times 10^{-12}"

"x=3.67\times 10^{-6}" "M"